OpenStudy (anonymous):

Solve for x: 2x - 4 = 6x + 8 pls help

OpenStudy (anonymous):

try to move all like terms to left side so first we move 6x to left hand side by subtracting 6x from both side 2x - 6x - 4 = 6x - 6x + 8 so we will get -4x - 4 = 8 now we get rid -4 from left side by adding 4 to both side -4x-4 + 4 = 8 + 4 so we will get -4x = 12 now to get rid of -4from left side we divide both sides by -4 so -4x/-4 = 12/-4 ==> x = -3